Circular No. 8778

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(42355), (58534), (65489) Further to IAUC 8756, K. S. Noll et al. report the discovery of a

companion to the distant object (65489) = 2003 FX_{128} (cf. MPEC 2003-H33, 2003-J40; MPS 78428) on images obtained with the Hubble Space Telescope on 2006 Apr. 11.916 UT; the companion at that time was ∼ 0.6 mag fainter than the primary and was located 0′′.085 ± 0′′.002 from the primary in p.a. 11^{o}.8 ± 0^{o}.2. The IAU’s Committee on Small-Body Nomenclature has approved the names Ceto for (65489) and Phorcys for the ‘binary’ companion (cf. MPC 57952). The CSBN has also approved the names Logos for the transneptunian object (58534) = 1997 CQ_{29} and Zoe for its ‘binary’ companion (cf. IAUC 7824, 7959; MPC 56962), as well as the names Typhon for (42355) = 2002 CR_{46} and Echidna for its ‘binary’ companion (cf. IAUC 8689; Noll et al. 2006, Icarus 184, 611; MPC 57951).

SUPERNOVAE 2006nr–2006oq Two apparent supernovae have been reported from unfiltered CCD im-

ages: 2006nr by T. Puckett and A. Pelloni (0.60-m reflector at Ellijay, GA; cf. IAUC 8771), and 2006oq by E. J. Christensen (Catalina Sky Survey).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006nr Oct. 26.24 5^{h}56^{m}23.^{s}51 +85^{o}54′38′′.8 15.8 41′′ E, 14′′ S
2006ob Nov. 18.29 1 51 48.11 + 0 15 48.3 17.6 5′′.5 W, 1′′.6 S
2006oq Nov. 17.49 11 44 00.07 +19 58 38.3 15.8 31′′ W, 17′′ N

Additional magnitudes for 2006nr in UGC 3336: Sept. 15, [19.5 (Puckett); Nov. 12.16, 16.2 (D. Lane, Stillwater Lake, NS, 0.28-m reflector). SN 2006nr is a type-Ia supernova ≈ 3 weeks past maximum on Nov. 18 (details on CBET 744). Nothing was visible at the position of 2006oq on 24 CSS images obtained between Mar. 2005 and Apr. 2006 (limiting mag V ∼ 21.5); type-Ia SN 2006oq was ∼ 2 weeks past maximum on Nov. 21 (details on CBET 748). On CBETs 743 and 745, J. Frieman and the Sloan Digital Sky Survey II collaboration report 24 additional supernovae (designated 2006ns–2006op, all in the magnitude range g = 19.8-22.3, and most being type-Ia events) discovered during Oct. 17–Nov. 17. The data in the table above for 2006ob are from M. Baek and W. Li (LOSS, unfiltered KAIT images); additional magnitudes for 2006ob: Oct. 29.30, [18.9 (LOSS); 7.31, [18.5 (LOSS); 13, 19.9 (SDSS II; g band); 19.31, 17.6 (LOSS). SN 2006ob was then an early type-Ia supernova. The non-type-Ia supernovae in this SDSS group are: 2006ns, II; 2006nx, Ib/c hypernova; 2006ny, IIp.

2006 November 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green